OpenStudy (anonymous):

Use elimination to solve. 5x-9y+53=0 9x-5y-53=0

OpenStudy (agreene):

\[(\frac{53}{4},\frac{53}{4})\]

OpenStudy (anonymous):

steps please.

OpenStudy (agreene):

5x-9y+53=0 x= 1/5 (-53+9 y) ------ 9x-5y-53=0 9(1/5 (-53+9 y))-5y-53=0 y= 53/4 ---- x= 1/5 (-53+9 y) x= 1/5 (-53+9 (53/4)) x= 53/4 (53/4,53/4) I dont think this is the method you asked for, but really this is the only method you need until you do complex variables... So, its all I remember :X